# Towards high-precision inspiral gravitational waveforms from binary neutron star mergers in numerical relativity

### v1

@@ I Introduction
NR is a chosen way to build a theoretical waveform template in the late inspiral phase of BNS mergers because (i) the NSs undergo significant tidal deformation and (ii) any analytic tools break down in this phase. Since 2010, the NR community started to make an effort to derive gravitational waveforms from BNS mergers whose quality should be fine enough to validate/calibrate the semi-analytical waveform models Kiuchi et al. (2018, 2020); Kawaguchi et al. (2018); Hotokezaka et al. (2013, 2015, 2016); Narikawa et al. (2019, 2020); Kuan et al. (2024, 2025); Dietrich et al. (2017a, 2018, 2019b, 2019a, b, c); Dietrich and Hinderer (2017); Bernuzzi et al. (2012a, b, 2014a, 2014b, 2015a); Foucart et al. (2019); Haas et al. (2016); Radice et al. (2014a, b); Baiotti et al. (2010).
@@
Among a couple of the error budgets in numerically derived gravitational waveforms, the truncation error due to the finite resolution is the most severe one. It implies that high-precision BNS merger simulations are necessary since numerical dissipation stemming from an approximate treatment of the partial differential equation, e.g., the finite difference, could deteriorate numerical solutions, resulting in a spurious acceleration of binary orbital decay. In other words, the gravitational wave phase error due to the numerical dissipation should be sufficiently suppressed compared to the gravitational wave phase shift due to the tidal deformation. Therefore, a higher-order scheme, particularly for a hydrodynamics solver, is an essential ingredient to achieve this requirement. To date, a couple of existing NR codes, Whicky-THC Radice et al. (2014a), BAM Bernuzzi and Dietrich (2016), FIL Most et al. (2019), and SpEC Foucart et al. (2021) implement a higher-order finite-difference scheme beyond the second order for relativistic (magneto)hydrodynamics to simulate BNS mergers (see SpECTRA Deppe et al. (2024) and Nmesh Tichy et al. (2023) for a discontinuous Galerkin method) (see also the waveform comparison among these NR codes in Ref. Hamilton and Messman (2025)).
@@
However, there is no NR code implementing a higher-order finite-volume method beyond the second order for relativistic hydrodynamics. Since the finite-volume method is mathematically more rigorous than the finite difference scheme to simulate astrophysical fluid dynamics in the sense that it relies on solving the so-called Riemann problem to capture the discontinuous solution, such as the shock wave, it is desired to implement a higher-order finite-volume Riemann solver in NR codes. Also, the quality of the Riemann solver should be improved. References Kiuchi et al. (2022); Kuan et al. (2024) demonstrate that the Harten-Lax-van Leer-contact (HLLC) Riemann solver Mignone and Bodo (2005) is superior to the Harten-Lax-van Leer-Einfeldt (HLLE) solver, which most existing NR codes implement. Particularly, not only the quantitative improvement of the non-spinning BNS merger dynamics Kiuchi et al. (2022) but also the qualitative improvement to simulate the $f$-mode resonance in the rapidly anti-spinning BNS merger are reported Kuan et al. (2024).
@@ II.1 Implementation
We implement a fourth-order HLLC Riemann solver in the framework of the finite-volume method. The readers may refer to Ref. Kiuchi et al. (2022) for the derivation of the finite-volume method and a tetrad transformation to a locally Minkowsky spacetime in a curved spacetime. In general, the higher-order Riemann solver beyond the second order is not straightforward in the finite-volume method because of the primitive-conservative conversion. We note that the higher-order finite difference Riemann solver is straightforward Radice et al. (2014a); Bernuzzi and Dietrich (2016); Most et al. (2019); Foucart et al. (2021).
@@
As a second step, we perform a primitive recovery for the fourth-order accurate point-wise conserved variable, $Q^{\text{PW}}$, to obtain the fourth-order accurate point-wise primitive variable, $P^{\text{PW}}$. Because the primitive recovery is done for the point-wise variable, it maintains the fourth-order accuracy.
@@
Practically, we implement the MP5 scheme Suresh and Huynh (1997) for the cell reconstruction, and the HLLC solver for the Riemann solver Mignone and Bodo (2005); Kiuchi et al. (2022). We also estimate the shock detector $\theta_{c}$ by using the five stencils as proposed in Ref. Berta et al. (2024) to distinguish the discontinuity and extrema (see their Eq. (36)). If the shock is detected, i.e., $\theta_{c}=0$ in a cell, we switch to a lower-order reconstruction scheme (the third-order accurate PPM scheme in our case) in this trouble cell.
@@
As a final step, we update the volume-averaged conserved variable with the numerical fluxes $\hat{f}_{j+1/2,k,l}$. In this way, the fourth-order accuracy is maintained in the finite-volume method.
@@ IV Discussion
The convergence property for the post-merger gravitational waveform is completely lost, particularly with the low resolution and/or the second-order HLLC (see the sign flip in the phase error in the top panel of Fig. 6). This behavior is observed in many other NR simulations for BNS mergers (see Ref. Dietrich et al. (2019a) for example). However, our new fourth-order HLLC could mitigate this pathological behavior in the post-merger gravitational waves if we employ a super-high resolution with $N\geq 189$ as indicated by (i) the increase of the density-weighted shock detector by improving the resolution during the post-merger phase in the left panel of Fig. 7 and (ii) no sign flip in the phase error between $N=189$ and $157$ during the post-merger phase (at least until the end of the simulation) in the top panel of Fig. 6.
@@ Appendix A Self-convergence test
where $N=109$ or $125$. $p$ denotes the convergence order. Figure 8 plots the phase error multiplied by the self-convergence factor with $p=2.5$ (the solid curves) estimated by the ”human eye” on top of the phase error between $N=109$ and $189$ (the green-dotted curve) and between $N=125$ and $189$ (the orange-dotted curve) derived by the second-order HLLC. Strictly speaking, the two curves, i.e., the original phase error and error multiplied by the self-convergence factor, do not perfectly overlap during the entire inspiral phase, which is also the case in the literature Dietrich et al. (2019a); Bernuzzi et al. (2015b). The color-shaded regions denote the phase error multiplied by the self-convergence factor, with the convergence order varying by $\pm 0.5$ from $2.5$, which also looks consistent with the original phase error. It demonstrates that the ”human eye’s inspection” of the convergence order in the self-convergence study is not stringent in the sense that it could overstimate/underestimate the convergence order.
Figure 8: The gravitational phase error as a function of the reterded time for the second-order HLLC. The green- and orange-dashed curves show the gravitational wave phase error between $N=109$ and $189$, and between $N=125$ and $189$, respectively. The green- and orange-solid curves denote the phase error between $N=157$ and $189$ multiplied by the self-convergence factor assuming $p=2.5$, respectively. The color shaded regions assume the convergence order with $2.5\pm 0.5$.

### v2

@@ I Introduction
NR is a chosen way to build a theoretical waveform template in the late inspiral phase of BNS mergers because (i) the NSs undergo significant tidal deformation and (ii) any analytic tools break down in this phase. Since 2010, the NR community started to make an effort to derive gravitational waveforms from BNS mergers whose quality should be fine enough to validate/calibrate the semi-analytical waveform models Kiuchi et al. (2017, 2020); Kawaguchi et al. (2018); Hotokezaka et al. (2013, 2015, 2016); Narikawa et al. (2019, 2020); Kuan et al. (2024, 2025); Dietrich et al. (2017a, 2018, 2019b, 2019a, b, c); Dietrich and Hinderer (2017); Bernuzzi et al. (2012a, b, 2014a, 2014b, 2015a); Foucart et al. (2019); Haas et al. (2016); Radice et al. (2014a, b); Baiotti et al. (2010); Gamba et al. (2021).
@@
Among a couple of the error budgets in numerically derived gravitational waveforms, the truncation error due to the finite resolution is the most severe one. It implies that high-precision BNS merger simulations are necessary since numerical dissipation stemming from an approximate treatment of the partial differential equation, e.g., the finite difference, could deteriorate numerical solutions, resulting in a spurious acceleration of binary orbital decay. In other words, the gravitational wave phase error due to the numerical dissipation should be sufficiently suppressed compared to the gravitational wave phase shift due to the tidal deformation. Therefore, a higher-order scheme, particularly for a hydrodynamics solver, is an essential ingredient to achieve this requirement. To date, a couple of existing NR codes, WhiskyTHC Radice et al. (2014a), BAM Bernuzzi and Dietrich (2016), FIL Most et al. (2019), and SpEC Foucart et al. (2021) implement a higher-order finite-difference scheme beyond the second order for relativistic (magneto)hydrodynamics to simulate BNS mergers (see SpECTRE Deppe et al. (2024) and Nmesh Tichy et al. (2023) for a discontinuous Galerkin method) (see also the waveform comparison among these NR codes in Ref. Hamilton and Messman (2025)).
@@
However, there is no NR code implementing a higher-order finite-volume method beyond the second order for relativistic hydrodynamics. Since the finite-volume method could be preferred over the finite difference scheme to simulate astrophysical fluid dynamics in the sense that it fully relies on solving the so-called Riemann problem to capture the discontinuous solution, such as the shock wave, it is desired to implement a higher-order finite-volume Riemann solver in NR codes. Also, the quality of the Riemann solver should be improved. References Kiuchi et al. (2022); Kuan et al. (2024) demonstrate that the Harten-Lax-van Leer-contact (HLLC) Riemann solver Mignone and Bodo (2005) is superior to the Harten-Lax-van Leer-Einfeldt (HLLE) solver, which most existing NR codes implement. Particularly, not only the quantitative improvement of the non-spinning BNS merger dynamics Kiuchi et al. (2022) but also the qualitative improvement to simulate the $f$-mode resonance in the rapidly anti-spinning BNS merger are reported Kuan et al. (2024).
@@ II.1 Implementation
We implement a fourth-order HLLC Riemann solver in the framework of the finite-volume method. The readers may refer to Ref. Kiuchi et al. (2022) for the derivation of the finite-volume method and a tetrad transformation to a locally Minkowski spacetime in a curved spacetime. In general, the higher-order Riemann solver beyond the second order is not straightforward in the finite-volume method because of the primitive-conservative conversion. We note that the higher-order finite difference Riemann solver is straightforward Radice et al. (2014a); Bernuzzi and Dietrich (2016); Most et al. (2019); Foucart et al. (2021).
@@
As a second step, we perform a primitive recovery for the fourth-order accurate point-wise conserved variable, $Q^{\text{PW}}$, to obtain the fourth-order accurate point-wise primitive variable, $P^{\text{PW}}$. Because the primitive recovery is done for the point-wise variable, it maintains the fourth-order accuracy. We should note that Eq. (2) cannot be applied to the point-wise primitive variable to maintain consistency between the point-wise conserved and primitive variables.
@@
Practically, we implement the MP5 scheme Suresh and Huynh (1997) for the cell reconstruction, and the HLLC solver for the Riemann solver Mignone and Bodo (2005); Kiuchi et al. (2022). We also estimate the shock detector $\theta_{c}$ by using the five stencils as proposed in Ref. Berta et al. (2024) to distinguish the discontinuity and extrema. Specifically, $\theta_{c}$ is estimated by
$\theta_{c}=\left\{\begin{array}[]{cc}1&\text{if }\eta_{c}<\eta_{d},\\ 0&\text{otherwise,}\end{array}\right$. (5)
where $\eta_{c}=\sqrt{\eta_{c,x}^{2}+\eta_{c,y}^{2}+\eta_{c,z}^{2}}$ and $\eta_{d}$ is a threshold parameter which is set to be $0.005$ in our implementation. For example, $\eta_{c,x}$ at $x_{j}$ is estimated by $\max(\eta^{o}_{c,x},\eta^{e}_{c,x})$ where
$\eta^{o}_{c,x}=\frac{|\delta^{(3)}\bar{Q}_{c}|}{|\bar{Q}_{\text{c,ref}}|+|\delta^{(1)}\bar{Q}_{c}|+|\delta^{(3)}\bar{Q}_{c}|+\epsilon},~\eta^{e}_{c,x}=\frac{|\delta^{(4)}\bar{Q}_{c}|}{|\bar{Q}_{\text{c,ref}}|+|\delta^{(2)}\bar{Q}_{c}|+|\delta^{(4)}\bar{Q}_{c}|+\epsilon},$ (6)
where
$\delta^{(1)}\bar{Q}_{c}=\frac{1}{2}\left(\bar{Q}_{j+1}-\bar{Q}_{j-1}\right),~\delta^{(2)}\bar{Q}_{c}=\left(\bar{Q}_{j+1}-2\bar{Q}_{j}+\bar{Q}_{j-1}\right),~\delta^{(3)}\bar{Q}_{c}=\frac{1}{2}\left(\bar{Q}_{j+2}-2\bar{Q}_{j+1}+2\bar{Q}_{j-1}-\bar{Q}_{j-2}\right),$
$\delta^{(4)}\bar{Q}_{c}=\bar{Q}_{j+2}-4\bar{Q}_{j+1}+6\bar{Q}_{j}-4\bar{Q}_{j-1}+\bar{Q}_{j-2},$ (7)
and we set $\epsilon=10^{-15}$ to avoid the denominator from being zero. We compute $\eta^{o/e}_{c,x}$ with the volume-averaged rest-mass density, pressure, and three velocities, respectively. For the reference value $\bar{Q}_{\text{c,ref}}$, we employ the volume-averaged rest-mass density and pressure, while we employ the iso-thermal sound speed $\sqrt{\bar{P}/\bar{\rho}}$ for the three velocity components.
If the shock is detected, i.e., $\theta_{c}=0$ in a cell, we switch to a lower-order reconstruction scheme (the third-order accurate PPM scheme in our case) in this trouble cell.
@@
As a final step, we update the volume-averaged conserved variable with the numerical fluxes $\hat{f}_{j+1/2,k,l}$ and execute the primitive recovery for the volume-averaged variables. In this way, the fourth-order accuracy is maintained in the finite-volume method.
@@
Before closing this subsection, we should mention the treatment of the source term in the Einstein equation and the hydrodynamics equation. We employ the point-wise variables for the source term in the Einstein equation. In a curved spacetime, the source term also appears in the relativistic Euler and energy equations. Since it is composed of a coupling of the geometrical and hydrodynamical variables, it is non-trivial to obtain the fourth-order accurate volume-averaged source term (see Eq. (3.15) in Ref. Kiuchi et al. (2022)). In this paper, we employ the volume-averaged variables for the source term in the hydrodynamics equation, assuming a constant profile for the determinant of the three metric and that for the spatial derivative of the other geometrical variables inside a cell. However, there is room for improvement or for constructing a true fourth-order source term.
@@ IV Discussion
The convergence property for the post-merger gravitational waveform is completely lost, particularly with the low resolution and/or the second-order HLLC (see the sign flip in the phase error in the top panel of Fig. 6). This behavior is observed in many other NR simulations for BNS mergers (see Ref. Dietrich et al. (2019a) for example). However, we speculate that our new fourth-order HLLC may mitigate this pathological behavior in the post-merger gravitational waves if we employ a super-high resolution with $N\geq 189$ as indicated by (i) the increase of the density-weighted shock detector by improving the resolution during the post-merger phase in the left panel of Fig. 7 and (ii) no sign flip in the phase error between $N=189$ and $157$ during the post-merger phase (at least until the end of the simulation) in the top panel of Fig. 6.
@@ Acknowledgements.
Data availability
The data that support the findings of this article are openly available SACRA-WAVEFORM-DATA-BANK .
@@ Appendix A Self-convergence test
where $N=109$ or $125$. $p$ denotes the convergence order. Figure 8 plots the phase error multiplied by the self-convergence factor with $p=2.5$ (the solid curves) estimated by the ”human eye” on top of the phase error between $N=109$ and $189$ (the green-dotted curve) and between $N=125$ and $189$ (the orange-dotted curve) derived by the second-order HLLC. Strictly speaking, the two curves, i.e., the original phase error and error multiplied by the self-convergence factor, do not perfectly overlap during the entire inspiral phase, which is also the case in the literature Dietrich et al. (2019a); Bernuzzi et al. (2015b). The color-shaded regions denote the phase error multiplied by the self-convergence factor, with the convergence order varying by $\pm 0.5$ from $2.5$, which also looks consistent with the original phase error. It demonstrates that the ”human eye’s inspection” of the convergence order in the self-convergence study is not stringent in the sense that it could overestimate/underestimate the convergence order.
Figure 8: The gravitational phase error as a function of the reterded time for the second-order HLLC. The green- and orange-dashed curves show the gravitational wave phase error between $N=109$ and $189$, and between $N=125$ and $189$, respectively. The green- and orange-solid curves denote the phase error between $N=157$ and $189$ multiplied by the self-convergence factor assuming $p=2.5$, respectively. The color-shaded regions assume the convergence order with $2.5\pm 0.5$.